# Efficient Methods for Non-stationary Online Learning

### v2

@@ Abstract
Non-stationary online learning has drawn much attention in recent years. In particular, dynamic regret and adaptive regret are proposed as two principled performance measures for online convex optimization in non-stationary environments. To optimize them, a two-layer online ensemble is usually deployed due to the inherent uncertainty of the non-stationarity, in which a group of base-learners are maintained and a meta-algorithm is employed to track the best one on the fly. However, the two-layer structure raises the concern about the computational complexity — those methods typically maintain $\mathcal{O}(\log T)$ base-learners simultaneously for a $T$-round online game and thus perform multiple projections onto the feasible domain per round, which becomes the computational bottleneck when the domain is complicated. In this paper, we first present efficient methods for optimizing dynamic regret and adaptive regret, which reduce the number of projections per round from $\mathcal{O}(\log T)$ to $1$. The obtained algorithms require only one gradient query and one function evaluation at each round. Our technique hinges on the reduction mechanism developed in parameter-free online learning and requires non-trivial twists on non-stationary online methods. Furthermore, we study an even strengthened measure, namely the “interval dynamic regret”, and reduce the number of projections per round from $\mathcal{O}(\log^{2}T)$ to $1$ to minimize it. Our reduction demonstrates great generalizability and can be applied to two important applications: online stochastic control and online principal component analysis, resulting in methods that are both efficient and optimal. Finally, empirical studies verify our theoretical findings.
@@ 1.1 Dynamic Regret and Adaptive Regret
[…]
@@
[…]
@@
It is worth noting that the relationship between dynamic regret and adaptive regret for OCO is generally unclear (Zhang, 2020, Section 5), even though a black-box reduction from dynamic regret to adaptive regret has been proved for the simper expert setting (i.e., online linear optimization over simplex) (Luo and Schapire, 2015, Theorem 4). Hence, the two measures are separately developed and many algorithms have been proposed, including algorithms for dynamic regret (Zinkevich, 2003; Hall and Willett, 2013; Zhang et al., 2018a; Zhao et al., 2020, 2024, 2021; Baby and Wang, 2021; Jacobsen and Cutkosky, 2022) and the ones for adaptive regret (Hazan and Seshadhri, 2009; Daniely et al., 2015; Jun et al., 2017; Zhang et al., 2018b, 2019). To achieve the best of both worlds, there are studies (Zhang et al., 2020; Cutkosky, 2020) optimizing both measures simultaneously by an even strengthened metric called interval dynamic regret, defined as
[…]
1.2 Two-layer Structure and Projection Complexity Issue
@@
The fundamental challenge of optimizing these two non-stationary regret measures is the uncertainty of the environmental non-stationarity. Concretely, to ensure the robustness to the unknown environments, dynamic regret aims to compete with any feasible comparator sequence, while adaptive regret examines the local performance over any intervals. The unknown comparators or unknown intervals bring considerable uncertainty to online optimization. To address the issue, a two-layer online ensemble structure is usually deployed to optimize the measures, where a set of base-learners are maintained to handle the different possibilities of online environments and a meta-algorithm is employed to combine them all and track the unknown best one. Such a meta-base framework successfully achieves many state-of-the-art results, including the $\mathcal{O}(\sqrt{T(1+P_{T})})$ dynamic regret (Zhang et al., 2018a) and the $\mathcal{O}(\sqrt{(F^{\mathbf{u}}_{T}+P_{T})(1+P_{T})})$ small-loss dynamic regret for smooth functions (Zhao et al., 2020), where $P_{T}=\sum_{t=2}^{T}\lVert\mathbf{u}_{t}-\mathbf{u}_{t-1}\rVert_{2}$ is the path length and $F^{\mathbf{u}}_{T}=\sum_{t=1}^{T}f_{t}(\mathbf{u}_{t})$ is the cumulative loss of comparators; as well as the $\mathcal{O}(\sqrt{\lvert I\rvert\log T})$ adaptive regret (Jun et al., 2017) and the $\mathcal{O}(\sqrt{F_{I}\log F_{I}\log F_{T}})$ small-loss adaptive regret for smooth functions (Zhang et al., 2019) for any interval $I=[r,s]\subseteq[T]$, where $F_{I}=\min_{\mathbf{x}\in\mathcal{X}}\sum_{t=r}^{s}f_{t}(\mathbf{x})$ and $F_{T}=\min_{\mathbf{x}\in\mathcal{X}}\sum_{t=1}^{T}f_{t}(\mathbf{x})$. Besides, an $\mathcal{O}(\sqrt{\lvert I\rvert(\log T+P_{I})})$ interval dynamic regret is also achieved by a two-layer (or even three-layer) structure (Zhang et al., 2020), where $P_{I}=\sum_{t=r}^{s}\lVert\mathbf{u}_{t}-\mathbf{u}_{t-1}\rVert_{2}$ is the path length over the interval.
@@
The two-layer methods have demonstrated great effectiveness in tackling non-stationary online environments, whereas the gain is at the price of heavier computations than the methods for minimizing static regret. While it is believed that additional computations are necessary for more robustness, we are wondering whether it is possible to pay for a “minimal” computation overhead for adapting to the non-stationarity. To this end, we focus on the popular first-order online methods and aim to streamline unnecessary computations while retaining the same regret guarantees. Arguably, the most computationally expensive step of each round is the projection onto the convex feasible domain, namely, the projection operation $\Pi_{\mathcal{X}}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\lVert\mathbf{x}-\mathbf{y}\rVert_{2}$ for a convex set $\mathcal{X}\subseteq\mathbb{R}^{d}$. Typical two-layer non-stationary online algorithms require maintaining $N=\mathcal{O}(\log T)$ base-learners simultaneously to cover the possibility of unknown environments. Define the projection complexity of online methods as the number of projections onto the feasible domain per round. Then, those non-stationary methods suffer an $\mathcal{O}(\log T)$ projection complexity, whereas standard online methods for static regret minimization require only $1$ projection onto the feasible domain per round such as online gradient descent (Zinkevich, 2003).
@@ 1.3 Our Contributions and Techniques
In this paper, we design a generic mechanism to reduce the projection complexity of many existing non-stationary methods from $\mathcal{O}(\log T)$ to $1$ without sacrificing the regret optimality, hence matching the projection complexity of stationary methods. Our reduction is inspired by the recent advance in parameter-free online learning (Cutkosky and Orabona, 2018; Mhammedi et al., 2019). The idea is simple: we reduce the original problem learned in the feasible domain $\mathcal{X}$ to an alternative one learned in a surrogate domain $\mathcal{Y}\supseteq\mathcal{X}$ such that the projection onto it is much cheaper, e.g., simply choosing $\mathcal{Y}$ as a properly scaled Euclidean ball; and moreover, a carefully designed surrogate loss is necessary for the alternative problem to retain the regret optimality. We reveal that a necessary condition for our reduction mechanism to deploy and reduce the projection complexity is that the non-stationary online algorithm shall query the function gradient once and evaluate the function value once per round. Several algorithms for the worst-case dynamic regret or adaptive regret already satisfy the requirements, so we can immediately deploy the reduction and obtain their efficient counterparts with the same regret guarantees and $1$ projection complexity. However, many non-stationary algorithms, particularly those designed for small-loss bounds, do not satisfy the requirement. Hence, we require non-trivial efforts to make them compatible. Due to this, we have developed a series of algorithms that achieve worst-case/small-loss dynamic regret and adaptive regret with one projection per round (actually, with one gradient query and one function evaluation per round as well).
@@
Despite that the reduction mechanism of this paper has been studied in parameter-free online learning, applying it to non-stationary online learning requires new ideas and non-trivial modifications. Here we highlight the technical innovation. The main challenge comes from the reduction condition mentioned earlier — as the surrogate loss involves the projection operation, our reduction requires the algorithm query one gradient and evaluate one function value at each round. However, many non-stationary algorithms do not satisfy the requirement, which is to be contrasted to the parameter-free algorithms such as MetaGrad (van Erven and Koolen, 2016) or its variants (Wang et al., 2019; Mhammedi et al., 2019) that naturally satisfy the condition. For example, the SACS algorithm (Zhang et al., 2019) enjoys the best known small-loss adaptive regret, yet the method requires $N$ gradient queries and $N+1$ function evaluations at each round, where $N=\mathcal{O}(\log T)$ is the number of base-learners. Thus, we have to dig into the algorithm and modify it to fit our reduction. First, we replace their meta-algorithm with Adapt-ML-Prod (Gaillard et al., 2014), an expert-tracking algorithm with a second-order regret with excess losses to accommodate the linearized loss that is used to ensure one gradient query per round. Second, we introduce a sequence of time-varying thresholds to adaptively determine the problem-dependent geometric covers in contrast to a fixed threshold used in their method. In particular, we register the cumulative loss of the final decisions rather than the base-learner’s one to compare it with the changing thresholds, which renders the design of one function value evaluation per round and also turns out to be crucial for achieving an improved small-loss bound that can recover the best known worst-case adaptive adaptive regret (by contrast, SACS cannot obtain optimal worst-case adaptive regret). Our algorithm only requires one projection/gradient query/function evaluation at each round, substantially improving the efficiency of SACS algorithm that requires $N$ projections/gradient queries/function evaluations per round.
@@
In this paper, extending our preliminary work (Zhao et al., 2022b), we further study a more strengthened measure, i.e., interval dynamic regret, and develop an efficient method to optimize it, reducing the number of projections from $\mathcal{O}(\log^{2}T)$ to $1$ using our reduction mechanism. Similar techniques developed for the adaptive regret are also required when analyzing the interval dynamic regret, and additional twists in the geometric covering intervals and base-algorithm are demanded due to the more complicated structure. Table 1 summarizes our main results in comparison with existing ones, in which we focus on the small-loss bounds for convex and smooth functions. Our reduction mechanism demonstrates great generalizability and can be applied to two important applications: (i) online non-stochastic control (Agarwal et al., 2019a), and (ii) online principal component analysis (Warmuth and Kuzmin, 2008). In both applications, the projection operation is complicated and time-consuming, which becomes even worse when leveraging these methods to handle non-stationary environments, since typically $\mathcal{O}(\log T)$ projections per round are required by following previous methods (Zhang et al., 2019; Zhao et al., 2023). We explore the structure of these two problems and propose efficient methods with $1$ projection per round to solve them under non-stationary environments, while ensuring optimal theoretical guarantees.
@@
D-Reg Zhao et al. (2020) $\mathcal{O}\left(\log T\right)$ $1$ $\mathcal{O}(\sqrt{(F^{\mathbf{u}}_{T}+P_{T})(1+P_{T})})$
Theorem 3 (ours) $1$ $1$ $\mathcal{O}(\sqrt{(F^{\mathbf{u}}_{T}+P_{T})(1+P_{T})})$
@@
Theorem 4 (ours) $1$ $1$ $1$ $\mathcal{O}(\min\{\sqrt{F_{I}\log F_{I}\log F_{T}},\sqrt{\lvert I\rvert\log T}\})$
Interval D-Reg Zhang et al. (2020) $\mathcal{O}(\log^{2}T)$ $\mathcal{O}(\log^{2}T)$ $\mathcal{O}(\log^{2}T)$ $\mathcal{O}(\sqrt{\lvert I\rvert(\log T+P_{I})})$
Theorem 5 (ours) $1$ $1$ $1$ $\mathcal{O}(\sqrt{(F^{\mathbf{u}}_{I}+P_{I})(\log(F_{I}^{\mathbf{u}}+P_{I})\cdot\log(F_{T}^{\mathbf{u}}+P_{T})+P_{I})})$
@@ 1.4 Assumptions
In fact, it is possible to relax Assumption 3 by requiring it to hold only over an extended domain $\mathcal{X}^{\prime}$, which is slightly larger than the original domain $\mathcal{X}$. Specifically, this extended domain is defined as $\mathcal{X}^{\prime}=\{\mathbf{x}+\mathbf{y}:\mathbf{x}\in\mathcal{X},\mathbf{y}\in\mathbb{R}^{d},\lVert\mathbf{y}\rVert_{2}\leq G/L\}$; in other words, $\mathcal{X}^{\prime}$ is the Minkowski sum of the original domain $\mathcal{X}$ and a small Euclidean ball.

### v3

@@ Abstract
Non-stationary online learning has drawn much attention in recent years. In particular, dynamic regret and adaptive regret are proposed as two principled performance measures for online convex optimization in non-stationary environments. To optimize them, a two-layer online ensemble is usually deployed due to the inherent uncertainty of non-stationarity, in which multiple base-learners are maintained and a meta-algorithm is employed to track the best one on the fly. However, the two-layer structure raises concerns about computational complexity — such methods typically maintain $\mathcal{O}(\log T)$ base-learners simultaneously for a $T$-round online game and thus perform multiple projections onto the feasible domain per round, which becomes the computational bottleneck when the domain is complicated. In this paper, we present efficient methods for optimizing dynamic regret and adaptive regret that reduce the number of projections per round from $\mathcal{O}(\log T)$ to $1$. The proposed algorithms require only one gradient query and one function evaluation at each round. Our technique hinges on the reduction mechanism developed in parameter-free online learning and requires non-trivial modifications for non-stationary online methods. Furthermore, we study an even stronger measure, namely “interval dynamic regret”, and reduce the number of projections per round from $\mathcal{O}(\log^{2}T)$ to $1$ for minimizing it. Our reduction demonstrates broad generality and applies to two important applications: online stochastic control and online principal component analysis, resulting in methods that are both efficient and optimal. Finally, empirical studies verify our theoretical findings.
@@ 1.1 Dynamic Regret and Adaptive Regret
$\textsc{D-Reg}_{T}(\mathbf{u}_{1},\ldots,\mathbf{u}_{T})=\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\sum_{t=1}^{T}f_{t}(\mathbf{u}_{t}).$ (1)
@@
$\textsc{A-Reg}_{T}(\lvert I\rvert)=\max_{[r,r+\tau-1]\subseteq[T]}\bigg{\{}\sum_{t=r}^{r+\tau-1}f_{t}(\mathbf{x}_{t})-\min_{\mathbf{u}\in\mathcal{X}}\sum_{t=r}^{r+\tau-1}f_{t}(\mathbf{u})\bigg{\}}.$ (2)
@@
Substantial research has been devoted to optimizing these two measures, including algorithms for dynamic regret (Zinkevich, 2003; Hall and Willett, 2013; György and Szepesvári, 2016; Zhang et al., 2018a; Zhao et al., 2020, 2021, 2024; Baby and Wang, 2021; Jacobsen and Cutkosky, 2022; Zhao et al., 2022a; Bai et al., 2022; Zhang et al., 2023; Li et al., 2024; Zhang et al., 2025) and adaptive regret (Hazan and Seshadhri, 2009; Daniely et al., 2015; Adamskiy et al., 2016; Jun et al., 2017; Zhang et al., 2018b, 2019; Cutkosky, 2020; Zhang et al., 2021; Yang et al., 2024). It is worth noting that the relationship between dynamic regret and adaptive regret in online convex optimization (OCO) remains largely unclear (Zhang, 2020, Section 5), even though a black-box reduction from dynamic regret to adaptive regret has been established for the simpler prediction-with-expert-advice setting (i.e., online linear optimization over simplex) (Luo and Schapire, 2015, Theorem 4). As a result, the two measures are typically developed independently.
To achieve the best of both worlds, several studies (Zhang et al., 2020; Cutkosky, 2020) have proposed optimizing both measures simultaneously by optimizing an even strengthened metric called interval dynamic regret, defined as
$\textsc{Interval-D-Reg}_{T}(\lvert I\rvert;\mathbf{u}_{1},\ldots,\mathbf{u}_{T})=\max_{[r,r+\tau-1]\subseteq[T]}\bigg{\{}\sum_{t=r}^{r+\tau-1}f_{t}(\mathbf{x}_{t})-\sum_{t=r}^{r+\tau-1}f_{t}(\mathbf{u}_{t})\bigg{\}}.$ (3)
This measure requires the online algorithm to compete with a sequence of time-varying comparators over any interval.
1.2 Two-layer Online Ensemble and Projection Complexity Issue
@@
The fundamental challenge of optimizing the above non-stationary regret measures is the uncertainty of environmental non-stationarity. Concretely, to ensure robustness in unknown environments, dynamic regret aims to compete with any feasible comparator sequence, while adaptive regret examines the local performance over any intervals; with interval dynamic regret is even more ambitious. The unknown comparators or/and unknown intervals bring considerable uncertainty to online optimization. To address the issue, a two-layer online ensemble structure is usually deployed to optimize the measures (Zhao et al., 2024), where a set of base-learners are maintained to handle the different possibilities of online environments and a meta-algorithm is employed to combine them all and track the unknown best one. Such a meta-base framework successfully achieves many state-of-the-art results, including the $\mathcal{O}(\sqrt{T(1+P_{T})})$ dynamic regret (Zhang et al., 2018a) and the $\mathcal{O}(\sqrt{(F^{\mathbf{u}}_{T}+P_{T})(1+P_{T})})$ small-loss dynamic regret for smooth functions (Zhao et al., 2020), where $P_{T}=\sum_{t=2}^{T}\lVert\mathbf{u}_{t}-\mathbf{u}_{t-1}\rVert_{2}$ is the path length and $F^{\mathbf{u}}_{T}=\sum_{t=1}^{T}f_{t}(\mathbf{u}_{t})$ is the cumulative loss of comparators; as well as the $\mathcal{O}(\sqrt{\lvert I\rvert\log T})$ adaptive regret (Jun et al., 2017) and the $\mathcal{O}(\sqrt{F_{I}\log F_{I}\log F_{T}})$ small-loss adaptive regret for smooth functions (Zhang et al., 2019) for any interval $I=[r,s]\subseteq[T]$, where $F_{I}=\min_{\mathbf{x}\in\mathcal{X}}\sum_{t=r}^{s}f_{t}(\mathbf{x})$ and $F_{T}=\min_{\mathbf{x}\in\mathcal{X}}\sum_{t=1}^{T}f_{t}(\mathbf{x})$. Besides, an $\mathcal{O}(\sqrt{\lvert I\rvert(\log T+P_{I})})$ interval dynamic regret is also achieved by a two-layer (or even three-layer) structure (Zhang et al., 2020), where $P_{I}=\sum_{t=r}^{s}\lVert\mathbf{u}_{t}-\mathbf{u}_{t-1}\rVert_{2}$ is the path length over the interval.
@@
The two-layer online ensemble methods have demonstrated great effectiveness in tackling non-stationary online environments, whereas the gain is at the price of heavier computations than the methods for minimizing static regret. While it is believed that additional computations are necessary for more robustness, we are wondering whether it is possible to pay for a “minimal” computation overhead for adapting to the non-stationarity. To this end, we focus on the popular first-order online methods and aim to streamline unnecessary computations while retaining the same regret guarantees. Arguably, the most computationally expensive step of each round is the projection onto the convex feasible domain, namely, the projection operation $\Pi_{\mathcal{X}}[\mathbf{y}]=\operatorname*{arg\,min}_{\mathbf{x}\in\mathcal{X}}\lVert\mathbf{x}-\mathbf{y}\rVert_{2}$ for a convex set $\mathcal{X}\subseteq\mathbb{R}^{d}$. Typical two-layer non-stationary online algorithms require maintaining $N=\mathcal{O}(\log T)$ base-learners simultaneously to cover the possibility of unknown environments. Define the projection complexity of online methods as the number of projections onto the feasible domain per round. Then, those non-stationary methods suffer an $\mathcal{O}(\log T)$ projection complexity, whereas standard online methods for static regret minimization require only $1$ projection onto the feasible domain per round such as online gradient descent (Zinkevich, 2003).
@@ 1.3 Our Contributions and Techniques
In this paper, we design a generic mechanism to reduce the projection complexity of many existing non-stationary methods from $\mathcal{O}(\log T)$ to $1$ without sacrificing regret optimality, thereby matching the projection complexity of stationary methods. Our reduction is inspired by recent advances in parameter-free online learning (Cutkosky and Orabona, 2018; Mhammedi et al., 2019). The idea is simple: we reduce the original problem learned in the feasible domain $\mathcal{X}$ to an alternative one learned in a surrogate domain $\mathcal{Y}\supseteq\mathcal{X}$ where projection is much cheaper, e.g., simply choosing $\mathcal{Y}$ as a properly scaled Euclidean ball; and moreover, a carefully designed surrogate loss is necessary for the alternative problem to preserve the regret optimality. We reveal that a necessary condition for our reduction mechanism to deploy and reduce the projection complexity is that the non-stationary online algorithm shall query the gradient once and evaluate the function value once per round. Several existing algorithms for obtaining the worst-case (problem-independent) dynamic regret or adaptive regret already satisfy the requirements, and our reduction immediately yields efficient counterparts with the same regret guarantees but only one projection per round. However, many non-stationary algorithms, particularly those designed for small-loss bounds, do not satisfy this condition, requiring non-trivial efforts to make them compatible. To this end, we develop a family of algorithms that achieve worst-case/small-loss dynamic regret and adaptive regret with with one projection, one gradient query, and one function evaluation per round.
@@
Despite that the reduction mechanism of this paper has been studied in parameter-free online learning, applying it to non-stationary online learning requires new ideas and non-trivial modifications. Here we highlight the technical innovation. The main challenge comes from the reduction condition mentioned earlier: since the surrogate loss involves the projection operation, our reduction requires the algorithm query one gradient and evaluate one function value at each round. However, many non-stationary algorithms do not satisfy the requirement, which is to be contrasted to the parameter-free algorithms such as MetaGrad (van Erven and Koolen, 2016) or its variants (Wang et al., 2019; Mhammedi et al., 2019) that naturally satisfy the condition. For example, the SACS algorithm (Zhang et al., 2019) enjoys the best known small-loss adaptive regret, yet the method requires $N$ gradient queries and $N+1$ function evaluations at each round, where $N=\mathcal{O}(\log T)$ is the number of base-learners. Thus, we have to dig into the algorithm and modify it to fit our reduction. First, we replace their meta-algorithm with Adapt-ML-Prod (Gaillard et al., 2014), an expert-tracking algorithm with a second-order regret with excess losses to accommodate the linearized loss that is used to ensure one gradient query per round. Second, we introduce a sequence of time-varying thresholds to adaptively determine the problem-dependent geometric covers in contrast to a fixed threshold used in their method. In particular, we register the cumulative loss of the final decisions rather than the base-learner’s one to compare it with the changing thresholds, which renders the design of one function value evaluation per round and also turns out to be crucial for achieving an improved small-loss bound that can recover the best known worst-case adaptive adaptive regret (by contrast, SACS cannot obtain optimal worst-case adaptive regret). Our algorithm only requires one projection/gradient query/function evaluation at each round, substantially improving the efficiency of SACS algorithm that requires $N$ projections/gradient queries/function evaluations per round.
@@
In this paper, extending our preliminary conference version (Zhao et al., 2022c), we further study a stronger measure, interval dynamic regret, and develop an efficient method that reduces its projection complexity from $\mathcal{O}(\log^{2}T)$ to $1$ using our reduction mechanism. Analyzing interval dynamic regret requires techniques similar to those for adaptive regret, along with additional refinements in the geometric covering intervals and base-algorithm to handle the more complex structure. Table 1 summarizes our main results in comparison with existing ones, in which we focus on the small-loss bounds for convex and smooth functions. Our reduction mechanism demonstrates great generality and can be applied to two important applications: (i) online non-stochastic control (Agarwal et al., 2019a), and (ii) online principal component analysis (Warmuth and Kuzmin, 2008). In both applications, the projection operation is complicated and time-consuming, which becomes even worse when leveraging these methods to handle non-stationary environments, since typically $\mathcal{O}(\log T)$ projections per round are required by following previous methods (Zhang et al., 2019; Zhao et al., 2023). We explore the structure of these two problems and propose efficient methods with $1$ projection per round to solve them under non-stationary environments, while ensuring optimal theoretical guarantees.
@@
D-Reg Zhao et al. (2024) $\mathcal{O}\left(\log T\right)$ 11 $\mathcal{O}(\sqrt{(F^{\mathbf{u}}_{T}+P_{T})(1+P_{T})})$
Theorem 3 (ours) 11 11 $\mathcal{O}(\sqrt{(F^{\mathbf{u}}_{T}+P_{T})(1+P_{T})})$
@@
Theorem 4 (ours) 11 11 11 $\mathcal{O}(\min\{\sqrt{F_{I}\log F_{I}\log F_{T}},\sqrt{\lvert I\rvert\log T}\})$
Interval
D-Reg
Zhang et al. (2020) $\mathcal{O}(\log^{2}T)$ $\mathcal{O}(\log^{2}T)$ $\mathcal{O}(\log^{2}T)$ $\mathcal{O}(\sqrt{\lvert I\rvert(\log T+P_{I})})$
Theorem 5 (ours) 11 11 11 $\mathcal{O}(\sqrt{(F^{\mathbf{u}}_{I}+P_{I})(\log(F_{I}^{\mathbf{u}}+P_{I})\cdot\log(F_{T}^{\mathbf{u}}+P_{T})+P_{I})})$
@@ 1.4 Assumptions
In fact, it is possible to relax Assumption 3 by requiring it to hold only over an extended domain $\mathcal{X}^{\prime}$, which is slightly larger than the original domain $\mathcal{X}$. Specifically, this extended domain is defined as $\mathcal{X}^{\prime}=\{\mathbf{x}+\mathbf{y}\mid\mathbf{x}\in\mathcal{X},\mathbf{y}\in\mathbb{R}^{d},\lVert\mathbf{y}\rVert_{2}\leq G/L\}$; in other words, $\mathcal{X}^{\prime}$ is the Minkowski sum of the original domain $\mathcal{X}$ and a small Euclidean ball.